# Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs): learning neural networks for designing neutral inclusions

### v1

@@ 3 The proposed method: CoCo-PINNs
This section introduces the CoCo-PINNs, their advantages, and the mathematical reasoning behind why neutral inclusions designed by training remain effective even in untrained background fields. We begin with the loss design corresponding to the imperfect interface problem eq. 2.1, whose solution exhibits a discontinuity across $\partial\Omega$ due to the imposed boundary conditions. We denote the solutions inside and outside $\Omega$ as $u^{\text{int}}$ and $u^{\text{ext}}$, respectively, and represent their neural networks’ approximations as $u_{\text{NN}}^{\text{int}}$ and $u_{\text{NN}}^{\text{ext}}$. We named these solutions as trained forward solutions. We aim to train the interface function, represented by $p^{(n)}$ for the truncated series approximation and $p_{\text{NN}}$ for the fully connected neural network approximation. The method utilizing $p^{(n)}$ is referred to as CoCo-PINNs, while the approach using $p_{\text{NN}}$ to represent the interface function is called classical PINNs.
@@ 3.1 Model design for the forward-inverse problem
We utilize three sets of collocation points: $\Omega^{\text{int}},\leavevmode\nobreak\ \Omega^{\text{ext}}$, and $\partial\Omega$, which are finite sets of points corresponding to the interior, exterior, and boundary of $\Omega$, respectively, with a slight abuse of notation for $\partial\Omega$. We select collocation points based on conformal mapping theory to handle PINNs in arbitrarily shaped domains, and provide a detailed methodology for this selection in Section D.1. To address the boundary conditions for $z\in\partial\Omega$, we use $x_{z}=z+\delta N$ and $y_{z}=z-\delta N$ for the limit of the boundary from the exterior and interior, respectively, with small $\delta>0$, and unit normal vector $N$.
@@
$\mathcal{L}_{\text{PDE}}^{\text{int}}$ $=\nabla\cdot\sigma_{c}\nabla u^{\text{int}}_{\text{NN}}$ for $\mbox{for }x\in\Omega^{\text{int}},$ (3.1)
$\mathcal{L}_{\text{PDE}}^{\text{ext}}$ $=\nabla\cdot\sigma_{m}\nabla u^{\text{ext}}_{\text{NN}}$ for $\mbox{for }x\in\Omega^{\text{ext}},$ (3.2)
$\mathcal{L}_{\text{bd}}^{(1)}$ $=p^{(n)}\left(u^{\text{ext}}_{\text{NN}}-u^{\text{int}}_{\text{NN}}\right)-\sigma_{m}\frac{\partial u^{\text{ext}}_{\text{NN}}}{\partial\nu}(x_{z})$ for $\mbox{for }z\in\partial\Omega,$ (3.3)
$\mathcal{L}_{\text{bd}}^{(2)}$ $=\sigma_{m}\frac{\partial u^{\text{ext}}_{\text{NN}}}{\partial\nu}(x_{z})-\sigma_{c}\frac{\partial u^{\text{int}}_{\text{NN}}}{\partial\nu}(y_{z})$ for $\mbox{for }z\in\partial\Omega,$ (3.4)
$\mathcal{L}_{\text{Neutral}}$ $=u^{\text{ext}}_{\text{NN}}-H$ for $\mbox{for }x\in\Omega^{\text{ext}}.$ (3.5)
with $\partial u/\partial\nu=\langle\nabla u,N\rangle$. In the case where we train using classical PINNs, we replace the interface function $p^{(n)}$ with $p_{\text{NN}}$. To enforce the non-negativity of the interface function, we introduce an additional loss function $\mathcal{L}_{\text{plus}}=\max\{0,-p\}$.
By combining all the loss functions with weight variables $\{w_{i}\}_{i=1}^{5}$, we define the total loss by
$\mathcal{L}_{\text{Total}}=\ &\frac{w_{1}}{|\Omega^{\text{int}}|}\sum_{x\in\Omega^{\text{int}}}\left(\mathcal{L}_{\text{PDE}}^{\text{int}}\right)^{2}+\frac{w_{1}}{|\Omega^{\text{ext}}|}\sum_{x\in\Omega^{\text{ext}}}\left(\mathcal{L}_{\text{PDE}}^{\text{ext}}\right)^{2}+\frac{w_{2}}{|\partial\Omega|}\sum_{z\in\partial\Omega}\big{(}\mathcal{L}_{\text{bd}}^{(1)}\big{)}^{2}\\ \quad+&\frac{w_{3}}{|\partial\Omega|}\sum_{z\in\partial\Omega}\big{(}\mathcal{L}_{\text{bd}}^{(2)}\big{)}^{2}+\frac{w_{4}}{|\Omega^{\text{ext}}|}\sum_{x\in\Omega^{\text{ext}}}\left(\mathcal{L}_{\text{Neutral}}\right)^{2}+w_{5}\sum_{x\in\partial\Omega}\mathcal{L}_{\text{plus}}.\end{aligned}$ (3.6)
Here, $|A|$ denotes the number of the elements in the set $A$. We then consider the following loss with regularization term:
$\mathcal{L}_{\text{Reg}}=\begin{cases}\displaystyle\mathcal{L}_{\text{total}}+\epsilon\Big{(}2\pi\gamma^{2}|p_{0}|^{2}+4\pi\gamma^{2}\textstyle\sum_{k=1}^{n}(1+k^{2})|p_{k}|^{2}\Big{)},&p=p^{(n)},\\ \displaystyle\mathcal{L}_{\text{total}}+\epsilon\|\mathbf{w}_{p}\|_{F}^{2},&p=p_{\text{NN}},\end{cases}$ (3.7)
@@
This type of regularization is commonly used to address ill-posed problems. We used the loss in eq. 3.7 for all experiments.
@@ 3.2 Neutral inclusion effects for untrained linear fields
In this section, we briefly explain the reason that the CoCo-PINNs can yield the neutral inclusion effect for untrained background fields. Since the governing equation eq. 2.1 is linear with respect to $H$, the following trivially holds by the properties of linear PDEs:
@@
By Theorem 3.1, one can expect to find a function $p(x)$ such that $\Omega$ is neutral to all linear fields $H$ by training with only two background fields, assuming such a $p(x)$ exists. Although the existence of this function has not yet been theoretically verified, experiments in this paper with various shapes demonstrate that, for given $\Omega$, there exists a $p(x)$ that produces the neutral inclusion effect, meaning that the perturbation $u-H$ is negligible for all directions of $H$.
@@
In all examples, we successfully identified a $p(x)$ with the neutral inclusion effect by training with only a single $H$. These results can be explained by the following theorem.
Theorem 3.2.
Let $\Omega$ with an interface function $p(x)$ be neutral for a single background field $H$. If the first rows of $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ given in eq. 2.4 are linearly independent, then $\Omega$ is neutral also for all linear fields $H(x)=ax_{1}+bx_{2}$ of arbitrary directions $(a,b)\in\mathbb{R}^{2}$.
The proof of Theorem 3.2 can be found in Appendix C. As a future direction, it would be interesting to either prove that the first rows of $\widetilde{A}_{1}$ and $\widetilde{A}_{2}$ are linearly independent for any $\Omega$, or to find counterexamples–namely, inclusions that are neutral in only one direction.
Remark 2.
@@
According to the universal approximation theorem, for a given interface function $p(x)$, the analytic solution $u_{p}$ on a bounded set and $p$ on $\partial\Omega$ can be approximated by neural networks. Additionally, by the Fourier analysis, $p(x)$ can be approximated by a truncated Fourier series $p^{(n)}$. In light of Theorem 3.2 and remark 1, we train using only a single linear field $H$.
${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p^{(n)}}}$${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{\text{NN}}}}$${p}$${\{p_{k}\}_{k=0}^{n}}$${u_{p}}$${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u_{\text{NN}}}}$niv.FourierThm. 2.1Cred.
Figure 3.1: Credibility scheme. $u_{\text{NN}},p_{\text{NN}},p^{(n)}$ are trained results by PINNs.
@@
Figure 3.1 demonstrates the operational principles of the neural networks we designed. We use ‘Univ.’ to represent the universal approximation theorem, ‘Fourier’ for the Fourier series expansion, Thm. 2.1 for the analytical solution derived from the mathematical result, and ‘Cred.’ for credibility. We note that $u_{p}$ is the analytic solution to eq. 2.1 associated with $p$ obtained either from $p^{(n)}$ via CoCo-PINNs or from $p_{\text{NN}}$ via classical PINNs. The credibility of $u_{p}$ is determined by whether $\|u_{p}-u_{\text{NN}}\|$ is small. Credibility is a crucial factor in the proposed PINNs’ schemes for identifying neutral inclusions with imperfect boundary conditions. If the trained forward solution $u_{\text{NN}}$, obtained alongside with $p^{(n)}$ or $p_{\text{NN}}$, is close to $u_{p}$, we can conclude that the neural networks have successfully identified the interface function, ensuring that $\Omega$ exhibits the neutral inclusion effect. This is because $(u_{\text{NN}}-H)$ has small values in $\Omega^{\text{ext}}$ by the definition of the loss function $\mathcal{L}_{\text{Neutral}}$. In other words, if the interface function provided by neural networks makes both $\|u_{p}-u_{\text{NN}}\|$ and $\mathcal{L}_{\text{Neutral}}$ small, then this interface function is the desired one. However, if $\|u_{p}-u_{\text{NN}}\|$ is not small, it becomes unclear whether the neural networks have failed to solve the inverse problem or the forward problem.
@@ 4 Experiments
We present the successful outcomes for designing the neutral inclusions by using the CoCo-PINNs, as well as the experiment results to verify the “reliability” in terms of credibility, consistency, and stability. Credibility indicates whether the trained forward solution closely approximates the analytic solution, which we assess by comparing the trained forward solution with the analytic solution derived in Choi & Lim (2024). Consistency focuses on whether the interface functions obtained from the classical PINNs and the CoCo-PINNs converge to the same result for re-experiments under identical environments. It’s worth noting that even if the neural networks succeed in fitting the forward solution and identifying the interface parameter in a specific experiment, this success may only occur occasionally. Consistency is aimed at determining whether the training outcomes are steady or merely the result of chance, and it can be utilized as an indicator of the steadiness of the model. training model. Lastly, stability refers to the sensitivity of a training model, examining how the model’s output changes in response to variations in environments of PDEs.
The inclusions with the shapes illustrated in Figure 4.1 will be used throughout this paper. Each shaped inclusion is defined by the conformal map given by eqs. D.1, D.2, D.3 and D.4 in Section D.3. We named the shapes of the inclusions ‘square’, ‘fish’, ‘kite’, and ‘spike’.
Figure 4.1: The shaped inclusions: square, fish, kite, and spike.

### v2

@@ 3 The proposed method: CoCo-PINNs
This section introduces the CoCo-PINNs, their advantages, and the mathematical reasoning behind why neutral inclusions designed by training remain effective even in untrained background fields. We begin with the loss design corresponding to the imperfect interface problem 1, whose solution exhibits a discontinuity across $\partial\Omega$ due to the imposed boundary conditions. We denote the solutions inside and outside $\Omega$ as $u^{\text{int}}$ and $u^{\text{ext}}$, respectively, and represent their neural networks’ approximations as $u_{\text{NN}}^{\text{int}}$ and $u_{\text{NN}}^{\text{ext}}$. We named these solutions as trained forward solutions. We aim to train the interface function, represented by $p^{(n)}$ for the truncated series approximation and $p_{\text{NN}}$ for the fully connected neural network approximation. The method utilizing $p^{(n)}$ is referred to as CoCo-PINNs, while the approach using $p_{\text{NN}}$ to represent the interface function is called classical PINNs.
@@ 3.1 Model design for the forward-inverse problem
We utilize three sets of collocation points: $\Omega^{\text{int}},~\Omega^{\text{ext}}$, and $\partial\Omega$, which are finite sets of points corresponding to the interior, exterior, and boundary of $\Omega$, respectively, with a slight abuse of notation for $\partial\Omega$. We select collocation points based on conformal mapping theory to handle PINNs in arbitrarily shaped domains, and provide a detailed methodology for this selection in Section 4.1.1. To address the boundary conditions for $z\in\partial\Omega$, we use $x_{z}=z+\delta N$ and $y_{z}=z-\delta N$ for the limit of the boundary from the exterior and interior, respectively, with small $\delta>0$, and unit normal vector $N$.
@@
$\mathcal{L}_{\text{PDE},u}^{\text{int}}$ $=\nabla\cdot\sigma_{c}\nabla u^{\text{int}}_{\text{NN}}$ $\mbox{for }x\in\Omega^{\text{int}},$ (3)
$\mathcal{L}_{\text{PDE},u}^{\text{ext}}$ $=\nabla\cdot\sigma_{m}\nabla u^{\text{ext}}_{\text{NN}}$ $\mbox{for }x\in\Omega^{\text{ext}},$ (4)
$\mathcal{L}_{\text{bd},u}^{(1)}$ $=p^{(n)}(z)\left(u^{\text{ext}}_{\text{NN}}(x_{z})-u^{\text{int}}_{\text{NN}}(y_{z})\right)-\sigma_{m}\frac{\partial u^{\text{ext}}_{\text{NN}}}{\partial\nu}(x_{z})$ $\mbox{for }z\in\partial\Omega,$ (5)
$\mathcal{L}_{\text{bd},u}^{(2)}$ $=\sigma_{m}\frac{\partial u^{\text{ext}}_{\text{NN}}}{\partial\nu}(x_{z})-\sigma_{c}\frac{\partial u^{\text{int}}_{\text{NN}}}{\partial\nu}(y_{z})$ $\mbox{for }z\in\partial\Omega,$ (6)
$\mathcal{L}_{\text{Neutral},u}$ $=u^{\text{ext}}_{\text{NN}}-H_{u}$ $\mbox{for }x\in\Omega^{\text{ext}}.$ (7)
with $\partial u/\partial\nu=\langle\nabla u,N\rangle$. In the case where we train using classical PINNs, we replace the interface function $p^{(n)}$ with $p_{\text{NN}}$.
By combining all the loss functions with weight variables $\{w_{i}\}_{i=1}^{5}$, we define the total loss with $u$ by
$\begin{aligned} \mathcal{L}_{\text{Total},u}=\ &\frac{w_{1}}{|\Omega^{\text{int}}|}\sum_{x\in\Omega^{\text{int}}}\left(\mathcal{L}_{\text{PDE}}^{\text{int}}\right)^{2}+\frac{w_{1}}{|\Omega^{\text{ext}}|}\sum_{x\in\Omega^{\text{ext}}}\left(\mathcal{L}_{\text{PDE}}^{\text{ext}}\right)^{2}\\ \qquad+&\frac{w_{2}}{|\partial\Omega|}\sum_{z\in\partial\Omega}\big(\mathcal{L}_{\text{bd}}^{(1)}\big)^{2}+\frac{w_{3}}{|\partial\Omega|}\sum_{z\in\partial\Omega}\big(\mathcal{L}_{\text{bd}}^{(2)}\big)^{2}+\frac{w_{4}}{|\Omega^{\text{ext}}|}\sum_{x\in\Omega^{\text{ext}}}\left(\mathcal{L}_{\text{Neutral}}\right)^{2}.\end{aligned}$ (8)
Here, $|A|$ denotes the number of elements in the set $A$. To enforce the non-negativity of the interface function, we introduce an additional loss function $\mathcal{L}_{\text{plus}}=\max\{0,-p\}$.
Define 3, 4, 5, 6, 7 and 8 in the same manner, replacing $u$ with $v$ and $H_{u}$ with $H_{v}$. Now, define the total loss by
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{Total,u}}+\mathcal{L}_{\text{Total},v}+w_{5}\sum_{x\in\partial\Omega}\mathcal{L}_{\text{plus}}.$
We then consider the following loss with a regularization term:
$\mathcal{L}_{\text{Reg}}=\begin{cases}\displaystyle\mathcal{L}_{\text{total}}+\epsilon\Big(2\pi\gamma^{2}|p_{0}|^{2}+4\pi\gamma^{2}\textstyle\sum_{k=1}^{n}(1+k^{2})|p_{k}|^{2}\Big),&p=p^{(n)},\\ \displaystyle\mathcal{L}_{\text{total}}+\epsilon\|\mathbf{w}_{p}\|_{F}^{2},&p=p_{\text{NN}},\end{cases}$ (9)
@@
This type of regularization is commonly used to address ill-posed problems. We used the loss in 9 for all experiments.
@@ 3.2 Neutral inclusion effects for untrained linear fields
In this section, we briefly explain the reason that the CoCo-PINNs can yield the neutral inclusion effect for untrained background fields. Since the governing equation 1 is linear with respect to $H$, the following trivially holds by the properties of linear PDEs:
@@
By Theorem 3.1, one can expect to find a function $p(x)$ such that $\Omega$ is neutral to all linear fields $H$ by training with only two background fields, assuming such a $p(x)$ exists. Although the existence of this function has not yet been theoretically verified, experiments in this paper with various shapes demonstrate that, for given $\Omega$, there exists a $p(x)$ that produces the neutral inclusion effect, meaning that the perturbation uHu-H is negligible for all directions of $H$.
The existence of an interface parameter that induces perfect neutrality (specifically, identifying the class of inclusion shapes that allow for such existence) is a mathematically significant question that remains unresolved. Since the interface parameters possess infinitely many degrees of freedom, it may be possible to derive analytic relations between the conformal mapping coefficients and the coefficients of the interface parameter that cancel the expansion coefficients of the solution $u$ up to a prescribed order. Such relations would be an important step toward establishing an existence theory for interface parameters that realize neutrality. Our computational approach—providing numerous numerical examples—could offer valuable insights to aid future analytical investigations into the neutral inclusion problem with imperfect interface.
@@
According to the universal approximation theorem, for a given interface function $p(x)$, the analytic solution $u_{p}$ on a bounded set and $p$ on $\partial\Omega$ can be approximated by neural networks. Additionally, by the Fourier analysis, $p(x)$ can be approximated by a truncated Fourier series $p^{(n)}$. In light of Theorem 3.1, we train using two linear field $H_{u}=x_{1}$ and $H_{v}=x_{2}$.
[…]
Figure 3: Credibility scheme: The solution $u_{\text{NN}}$ is obtained from either classical PINNs or CoCo-PINNs. The interface function $p_{\text{NN}}$ is identified via classical PINNs, while the truncated series representation $p^{(n)}$ is inferred through the proposed CoCo-PINNs framework. This scheme enables a direct comparison between the PINN-predicted forward solutions and the analytic solution derived from reconstructed interface parameters, thereby providing a basis for evaluating the credibility of each method.
@@
Figure 3 demonstrates the operational principles of the neural networks we designed. We use ‘Univ.’ to represent the universal approximation theorem, ‘Fourier’ for the Fourier series expansion, [45] for the analytical solution derived from the mathematical result, and ‘Cred.’ for credibility. We note that $u_{p}$ is the analytic solution to 1 associated with $p$ obtained either from $p^{(n)}$ via CoCo-PINNs or from $p_{\text{NN}}$ via classical PINNs. The credibility of $u_{p}$ is evaluated based on the closeness of $u_{\text{NN}}$ to $u_{p}$, measured by the two norms $\|\cdot\|$ introduced in 10 and 11. Credibility is a crucial factor in the proposed PINNs’ schemes for identifying neutral inclusions with imperfect boundary conditions. If the trained forward solution $u_{\text{NN}}$, obtained alongside with $p^{(n)}$ or $p_{\text{NN}}$, is close to $u_{p}$, we can conclude that the neural networks have successfully identified the interface function, ensuring that $\Omega$ exhibits the neutral inclusion effect. This is because $(u_{\text{NN}}-H)$ has small values in $\Omega^{\text{ext}}$ by the definition of the loss function $\mathcal{L}_{\text{Neutral}}$. In other words, if the interface function provided by the neural network leads to both a small discrepancy between $u_{\text{NN}}$ and $u_{p}$, and a small value of the neutral loss $\mathcal{L}_{\text{Neutral}}$, we can consider the interface function to be valid. However, if $|u_{p}-u_{\text{NN}}|$ remains large, it becomes unclear whether the discrepancy is due to a failure in reconstructing the interface function (inverse problem) or in predicting the forward solution. To make this reasoning more precise, observe that
$\|u_{p}-H_{u}\|\leq\|u_{p}-u_{\text{NN}}^{\text{ext}}\|+\|u_{\text{NN}}^{\text{ext}}-H_{u}\|=\text{Cred.}+\mathcal{L}_{\text{Neutral},u}.$
Here, the first term $\|u_{p}-u_{\text{NN}}^{\text{ext}}\|$ quantifies the credibility of the network’s forward prediction (denoted as Cred.), while the second term $\|u_{\text{NN}}^{\text{ext}}-H_{u}\|$ corresponds to the neutral loss $\mathcal{L}_{\text{Neutral},u}$. This inequality illustrates that both the credibility and the neutral loss jointly contribute to validating the reconstructed interface function.
@@ 4 Experiments
We present the successful outcomes for designing the neutral inclusions by using the CoCo-PINNs, as well as the experiment results to verify the “reliability” in terms of credibility, consistency, and stability. Credibility indicates whether the trained forward solution closely approximates the analytic solution, which we assess by comparing the trained forward solution with the analytic solution derived in [45]. Consistency focuses on whether the interface functions obtained from the classical PINNs and the CoCo-PINNs converge to the same result for re-experiments under identical environments. It’s worth noting that even if the neural networks succeed in fitting the forward solution and identifying the interface parameter in a specific experiment, this success may only occur occasionally. Consistency is aimed at determining whether the training outcomes are steady or merely the result of chance, and it can be utilized as an indicator of the steadiness of the training model. Lastly, stability refers to the sensitivity of a training model, examining how the model’s output changes in response to variations in environments of PDEs.
Figure 4: Inclusion shapes: Square, fish, kite, and spike geometries are selected to represent varying degrees of geometric complexity in the assessment.
The inclusions with the shapes illustrated in Figure 4 will be used throughout this paper. Each shaped inclusion is defined by the conformal map given by 13, 14, 15 and 16 in Section 4.1.3. We named the shapes of the inclusions ‘square’, ‘fish’, ‘kite’, and ‘spike’.